how many full cycles of the function y=2cos3x appear in 2pi radians?
(1) 1
(2) 2
(3) 3
(4) 3pi/2

Answer by lwsshak3(11628)

How many full cycles of the function y=2cos3x appear in 2pi radians?
Equation for cos function: y=Acos(Bx-C), A=amplitude, period=2π/B, phase shift=C/B
For given problem: y=2cos3x
B=3
period=2π/B=2π/3
Number of full cycles in 2pi radians=2π/(2π/3)=3